Session: June 2008 (session 2)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (A. Cattani at Sun Jun 1 09:40:02 2008) Success :

Onsala (Michael Lindqvist at Sun Jun 1 09:44:49 2008) Minor failures :
Problem with "noisy" RCP channel back (see em059c). Have not been able to find any problems with the receiver (yet).

Effelsberg (Uwe Bach at Mon Jun 2 08:43:34 2008) Minor failures : POIN
Requested was EB_VLBA, but VLBA VME System was not working. Changed schedule to MKIV. Some uncertainties in the pointing were observed.

Effelsberg (Uwe Bach at Mon Jun 2 11:03:14 2008) Minor failures : POIN
More precisely: Pointing degrades between 04:00-06:30 (following sunrise). Point err of 11-15 arcsec corrected 06:28. Lost scans at 06:28 and 07:51 for pointing.

Robledo (Esther Moll at Thu Jun 5 13:31:14 2008) Minor failures : RECE POIN
1. Formatter was on sync. 2. Antenna problems: ONSOURCE info is not reliable (xant version 5.3 does not work properly). Autocollimator going in and out of lock for high elevation sources. As a result, there was bad antenna pointing for sources with elevation 79 deg or higher. 3. Subreflector properly tracking. 4. Mark5 recorder#1 performance OK. 1138390940712 bytes of data recorded. All scan_check correct. 5. Video converters#1 and 2 levels were very low because its frequecy setup (22196.49 MHZ 8MHz ) were out of the preselector band (2224535MHZ for MMS#1 and 2223736MHZ for MMS#2 ). RCP video converters: Signal level was correct for the rest of RCP video converters: VC#4, 6 and 8 (those connected to MMS#2). LCP video converters: MMS#1 (LCP) signal was not stable. The ifd value was adjusted several times between sources. Video converters# 3, 5 and 7 affected (VC#1 was out of preselector band). 6. Phase cal tones not currently available for K-Band. 7. Antabfs file generated. Experiment log and antabfs file transferred to vlbeer. 8. gps-fmout provided at experiment log.

Noto (P. Cassaro at Tue Jun 10 10:12:51 2008) Success :
No known problems

Jodrell_Bank (Alastair Gunn at Tue Jun 10 23:47:24 2008) Minor failures : ANT
The MkII telescope was used (Jb2). Telescope motors dropped out briefly at 05:51 UT, back on source within 1 minute. No other known problems.

Cambridge (Alastair Gunn at Tue Jun 10 23:47:54 2008) Success :
No known problems.